Two varieties of wheat are being tested in a developing country. Twelve test plots are given identical preparatory treatment. Six plots are sown with Variety 1 and the other six plots with Variety 2 in an experiment in which the crop scientist hope to determine whether there is significant difference between yields. Consider the following sample statistics: . What is the standard deviation of ? Assume equal variance.


1
Expert's answer
2021-09-09T16:00:02-0400


If we knew the real distribution average, we could construct a sum of squares of independent identically distributed standard normal variables "\u03be_i" similar to "S^2". It would've been "χ^2_n -distributed: \frac{(n−1)S^2}{σ^2}=∑_i^n(\frac{ξi−μ}{σ})^2∼χ^2_n" with n degrees of freedom.

However, we need to know the real distribution mean μ and distribution variance σ2 to achieve the desired sum, and they are not available in real-life small sample situations.

As we don't know the real distribution mean μ, you have to approximate it with sample average.

"ξ¯=\frac{∑_1^nξi}{n}"

But this takes away one degree of freedom (if you know the sample mean, then only ξi from 1 to n−1 can take arbitrary values, but the nth has to be "\u03be_n=\u03be\u2212\u2211_1^n\u22121\u03be_i )."

So our real S2 loses one degree of freedom "\frac{(n−1)S^2}{σ^2}=∑_1^n(\frac{ξi−ξ}{σ})^2∼χ^2_{n−1}"